跳到论文正文
W Papers独立论文阅读器 arXiv:1804.04948 论文正文 计算阅读时间 · 0%
引用这篇论文根据当前论文元数据生成,正式投稿前建议核对。

    
  
License: arXiv.org perpetual non-exclusive license
arXiv:1804.04948v4 [stat.OT] 22 Oct 2018

The rationality of irrationality in the Monty Hall problem

Torsten Enßlin    Margret Westerkamp Affiliation: Max Planck Institute for Astrophysics, Karl-Schwarzschild-Str. 1, 85741 Garching, Germany
Abstract

The rational solution of the Monty Hall problem unsettles many people. Most people, including the authors, think it feels wrong to switch the initial choice of one of the three doors, despite having fully accepted the mathematical proof for its superiority. Many people, if given the choice to switch, think the chances are fifty-fifty between their options, but still strongly prefer to stay with their initial choice. Is there some sense behind these irrational feelings?

We entertain the possibility that intuition solves the problem of how to behave in a real game show, not in the abstract textbook version of the Monty Hall problem. A real showmaster sometimes plays evil, either to make the show more interesting, to save money, or because he is in a bad mood. A moody showmaster erases any information advantage the guest could extract by him opening other doors which drives the chance of the car being behind the chosen door towards fifty percent. Furthermore, the showmaster could try to read or manipulate the guest’s strategy to the guest’s disadvantage. Given this, the preference to stay with the initial choice turns out to be a very rational defense strategy of the show’s guest against the threat of being manipulated by its host. Thus, the intuitive feelings most people have about the Monty Hall problem coincide with what would be a rational strategy for a real-world game show. Although these investigations are mainly intended to be an entertaining mathematical commentary on an information-theoretic puzzle, they touch on interesting psychological questions.

The Monty Hall problem

The textbook Monty Hall problem [1, 2]goes as follows. A showmaster presents you three doors to choose from. Behind one is a sports car that you win if you decide to open this door. Behind the others are goats that have no value to you. You indicate a choice, but before you open the chosen door, the showmaster opens another one, revealing a goat, and gives you the opportunity to revise your choice. Should you stay with your choice or should you switch to the other door?

The implicit assumption is that the showmaster always behaves in this way. In this case the optimal strategy can be found by various lines of argument. Your initial probability of picking the right door is 1/3\nicefrac{{1}}{{3}}. If your strategy is to always stay with your choice, this is your probability of winning the car, irrespective of whether the showmaster opens zero, one, or both other doors. In 2/3\nicefrac{{2}}{{3}} of the cases the car is behind one of the other doors. The showmaster is so friendly as to indicate which of those is certainly wrong. So, if you switch, and the car is behind one of the other doors, he guides you to the good one. This will work in 2/3\nicefrac{{2}}{{3}} of the cases, which is then the chance of winning if you switch.

So far so good. Logic has shown us what to do best. But – it just feels wrong. It feels so wrong that approximately 10,00010,000 readers of Marilyn vos Savant’s column “Ask Marilyn” write to her to say that she must be erring after she had explained the correct solution to the Monty Hall problem. Among them, nearly 1,0001,000 holders of PhDs [3]. Why don’t they (and we) trust Marilyn? Well, probably, because for this we also have to trust the showmaster!

The evil showmaster

Imagine the textbook assumption is wrong that the showmaster always behaves in the described way. Real showmasters can and occasionally do behave differently. Imagine the worst case: the showmaster is evil [4] and opens other doors only if you pick the door with the car. Thus, if you pick a door with a goat, he will immediately open it and you lose. This happens in 2/3\nicefrac{{2}}{{3}} of the cases. In 1/3\nicefrac{{1}}{{3}} of the cases you pick the door with the car. Then, the evil showmaster opens another one in the hope that you play switch as you have learned this to be superior. In that case you lose as well. So if you always play switch and the showmaster plays evil, you always lose. However, if you play stay, you win in 1/3\nicefrac{{1}}{{3}} of the cases, namely, whenever you initially picked the door with the car. The showmaster cannot change this. Just be stubborn and ignore the seduction of a virtual additional chance of 1/3\nicefrac{{1}}{{3}} by switching and you defend the guaranteed 1/3\nicefrac{{1}}{{3}} winning chance of your initial choice.

A showmaster always playing evil would save the broadcasting company real money. Only 1/3\nicefrac{{1}}{{3}} of the stay-people get cars, and all the switch-people get only goats. However, he might ruin the show. It would become boring and people would stop watching it. Therefore, he should behave in an evil way only with some frequency pp. And he should tailor this frequency to create maximal suspense while cutting financial losses.

Optimal moodiness

In order to optimally choose his frequency pp of playing evil the showmaster has to consider the situation of his guests. For this, he had better assume that whatever frequency pp he chooses, this will also be known to his guests. They might obtain this number by studying broadcasts of the show or by performing the same calculation as he does to determine his optimal moodiness, as we will do in the following as well.

Imagine that you have made your initial choice, and he has opened another door, revealing a goat. Should you now stay or switch? In the fraction pp of the cases when his mood is evil you had better stay, ensuring your winning probability of 1/3\nicefrac{{1}}{{3}}. If you switch in this case, you surely lose. In the fraction 1−p1-p of the cases, when he plays the fair Monty Hall textbook showmaster, you win in 1/3\nicefrac{{1}}{{3}} of the cases if you stay and in 2/3\nicefrac{{2}}{{3}} if you switch. Multiplying and adding these cases together yields that you win with playing staying in p×1/3+(1−p)×1/3=1/3p\times\nicefrac{{1}}{{3}}+(1-p)\times\nicefrac{{1}}{{3}}=\nicefrac{{1}}{{3}} of all cases. If you play switching, you win in p×0+(1−p)×2/3=(1−p)×2/3p\times 0+(1-p)\times\nicefrac{{2}}{{3}}=(1-p)\times\nicefrac{{2}}{{3}} of the cases. Which strategy is better depends on pp. If p<1/2p<\nicefrac{{1}}{{2}}, switching is better, if p>1/2p>\nicefrac{{1}}{{2}}, staying is better, and for p=1/2p=\nicefrac{{1}}{{2}} the strategy does not matter, you win in 1/3\nicefrac{{1}}{{3}} of all cases.

The showmaster should make you face a fifty-fifty chance of him being evil, as this will put you under maximal stress. This is what the audience wants to see and therefore this is what the showmaster will probably aim for. He might add or subtract a small margin, in order to save a bit money or to make the show more attractive, respectively, but this is hard for you to judge. He will make sure that you are nearly clueless about his mood by choosing p≈1/2p\approx\nicefrac{{1}}{{2}}. This optimal amount of evilness is therefore what you better assume. He will be moody just to the level that erases any information advantage you could have gotten from having observed his action.

The information game

The probability p=P⁡(evil)p=P(\mbox{evil}) describes your belief before the game that the showmaster has evil intentions. During the game the probability you assign to him being evil will change.

If he opens the door that you initially picked and reveals a goat there (“my” for “he opens my door”), then you can be sure that he is evil, P⁡(evil | my)=1P(\text{evil | my})=1, as the “fair” showmaster never does this, but always opens an “other” door, P⁡(my | fair)=1−P⁡(other | fair)=0P(\text{my | fair})=1-P(\text{other | fair})=0. We assume here that no showmaster ever reveals the sports car without you insisting on choosing its door, P⁡(other | car)=1P(\text{other | car})=1, where with “car” and “goat” we label what is behind your initially chosen door.

In the other case, where he opens another door, the situation is less clear. However, since some of the evil possibilities are ruled out, your confidence in him being in a fair mood has increased. How much can be worked out using Bayes’ theorem as we have our prior beliefs

P⁡(evil)=p​ and ​P​(fair)=1−pCLOSEP(\text{evil)}=p\mbox{ and }P(\text{fair})=1-p

and can specify the likelihoods of the different events:

P⁡(other | fair)\displaystyle P(\text{other | fair}) =1,\displaystyle=1, P⁡(my | fair)CLOSE\displaystyle P(\text{my | fair)} =0,\displaystyle=0,
P⁡(other | evil)\displaystyle P(\text{other | evil}) =13,\displaystyle=\frac{1}{3}, P⁡(my | evil)\displaystyle P(\text{my | evil}) =23.\displaystyle=\frac{2}{3}.

Going through the maths of Bayes’ theorem we find

P⁡(evil | other)\displaystyle P(\mbox{evil }|\mbox{ other}) =\displaystyle= P⁡(other, evil)P⁡(other)CLOSE\displaystyle\frac{P(\mbox{other, evil})}{P(\text{other)}}
=\displaystyle= P⁡(other | evil)​P​(evil)P⁡(other, evil)+P⁡(other, fair)\displaystyle\frac{P(\mbox{other }|\mbox{ evil})\,P(\text{evil})}{P(\text{other, evil})+P(\text{other, fair})}
=\displaystyle= 13×p13×p+1×(1−p)\displaystyle\frac{\frac{1}{3}\times p}{\frac{1}{3}\times p+1\times(1-p)}
=\displaystyle= p3−2​p.\displaystyle\frac{p}{3-2\,p}.

We have P⁡(evil | other)=1/4P(\mbox{evil }|\mbox{ other})=\nicefrac{{1}}{{4}} for p=1/2p=\nicefrac{{1}}{{2}}. Thus, having seen the showmaster opening another door makes you more confident in him being fair. Should you therefore trust him and open the remaining door as now your probability of him being evil is well below 1/2\nicefrac{{1}}{{2}}?

Not at all! You should still play carefully, as the threshold of p=1/2p=\nicefrac{{1}}{{2}} was derived for the prior probability P⁡(evil)P(\mbox{evil}) and not for the posterior probability P⁡(evil | other)P(\mbox{evil }|\mbox{ other}). In fact all that matters to you is P⁡(car | other)CLOSEP(\mbox{car |\mbox{ other})}, the probability that the car is behind your initially picked door, irrespective of whether the showmaster is evil or not. This requires you to marginalize out the momentary mood of the showmaster, but not his action, as the latter is important information. For calculating this chance, we recall that P⁡(other | car)=1P(\text{other | car})=1 as any showmaster, regardless of fair or evil, will open another door if you picked the car. If you picked a goat, only the fair one will open another door. This will therefore happen with probability P⁡(other | goat)=1−pP(\text{other | goat})=1-p. The probabilities of seeing another door opened given what is behind your door now permit us to work out

P⁡(car | other)CLOSE\displaystyle P(\mbox{car |\mbox{ other})} =\displaystyle= P⁡(other, car)P⁡(other)CLOSE\displaystyle\frac{P(\mbox{other, car})}{P(\mbox{other)}}
=\displaystyle= P⁡(other | car)​P​(car)P⁡(other, car)+P⁡(other, goat)\displaystyle\frac{P(\text{other | car})\,P(\text{car})}{P(\mbox{other, car})+P(\mbox{other, goat})}
=\displaystyle= 1×131×13+(1−p)×23\displaystyle\frac{1\times\frac{1}{3}}{1\times\frac{1}{3}+(1-p)\times\frac{2}{3}}
=\displaystyle= 13−2​p.\displaystyle\frac{1}{3-2p}.

Thus, a showmaster with evil intentions half of the time leaves us clueless behind which of the remaining doors the car is to be expected, since P⁡(car | other)=1/2CLOSEP(\mbox{car |\mbox{ other})}=\nicefrac{{1}}{{2}} for p=1/2p=\nicefrac{{1}}{{2}}. This value is the intuitive feeling of many people about their chance, indicating that they also anticipate a fifty-fifty percent chance of the showmaster playing evil, an option a billion years of evolution have taught us never to forget. Or at least, our feeling tells us that the showmaster is manipulating us and probably has erased any information we could have obtained on the correct door.

In contrast to this, a showmaster always exhibiting the same mood would provide you with information on the door with the car. His message, however, would depend on his mood. A showmaster always playing fairly (p=0p=0) would leave only a winning chance of P⁡(car | other, fair)=1/3CLOSEP(\mbox{car |\mbox{ other, fair})}=\nicefrac{{1}}{{3}} for your door, but 2/3\nicefrac{{2}}{{3}} for the remaining door. The always evil showmaster (p=1p=1) would inform us by opening another door that we picked the right one, as P⁡(car | other, evil)=1CLOSEP(\mbox{car |\mbox{ other, evil})}=1 then. To benefit from this message, you would need to have identified his mood. But beware, he might have read your mind as well.

The mind-reading showmaster

If a fraction q=P⁡(stay)q=P(\text{stay}) of the guests of the show stay with their doors and 1−q=P⁡(switch)1-q=P(\text{switch}) of the guests switch to the other door, the fraction of guests winning the car is

P⁡(win)\displaystyle P(\text{win}) =\displaystyle= P⁡(stay)​P​(car)+P⁡(switch)​P​(fair)​P​(goat)CLOSE\displaystyle P(\text{stay)}\,P(\text{car})+P(\text{switch})\,P(\text{fair})\,P(\text{goat})
=\displaystyle= q×13+(1−q)×(1−p)×23\displaystyle q\times\frac{1}{3}+(1-q)\times(1-p)\,\times\frac{2}{3}
=\displaystyle= 2−2​p−q+2​p​q3\displaystyle\frac{2-2\,p-q+2\,p\,q}{3}

as “stay” wins irrespective of the showmaster’s strategy if the car is behind the initial door, and "switch" wins if there was a goat and the showmaster plays fair. For the optimally moody showmaster with p=1/2p=\nicefrac{{1}}{{2}} this means P⁡(win)=1/3P(\text{win})=\nicefrac{{1}}{{3}} as the guest strategy then does not matter. The always fair showmaster is significantly more expensive for the broadcasting company, as for p=0p=0 we have P⁡(win)=(2−q)/3≥1/3P(\text{win})=\nicefrac{{(2-q)}}{{3}}\geq\nicefrac{{1}}{{3}} and thus the switching guests double their chance. The always evil showmaster (p=1p=1) saves real money with P⁡(win)=q/3≤1/3P(\text{win})=\nicefrac{{q}}{{3}}\leq\nicefrac{{1}}{{3}} as all switching guests lose surely. However, he can follow this strategy only until the reputation of the show is ruined.

Anyhow, such an economic winning rate could also be achieved by a mind-reading showmaster, who adapts his strategy to the situation. He could play fair whenever there is a guest expected to play stay or one who will switch to a goat. Otherwise he plays evil. Then the number of winning guests would be that of the always evil master. But the showmaster would look much more generous than the evil one, as he opens other doors more frequently.

If the showmaster can read his guest’s mind perfectly, the winning rate is P⁡(win)=q/3≤1/3P(\text{win})=\nicefrac{{q}}{{3}}\leq\nicefrac{{1}}{{3}}, but he opens other doors in

P⁡(other | mind-reader)\displaystyle P(\text{other | mind-reader}) =\displaystyle= P⁡(stay)+P⁡(switch)​P​(car)\displaystyle P(\text{stay})+P(\text{switch})\,P(\text{car})
=\displaystyle= q+(1−q)×13\displaystyle q+(1-q)\times\frac{1}{3}
=\displaystyle= 1+2​q3\displaystyle\frac{1+2\,q}{3}

of the games, more often than the 1/3\nicefrac{{1}}{{3}} if always playing evil. This will look generous, given that most people prefer to stay in fear of an evil showmaster, meaning q>1/2q>\nicefrac{{1}}{{2}}, which implies P⁡(other | mind reader)>2/3P(\text{other | mind reader})>\nicefrac{{2}}{{3}}. With a bit of mind-reading, the showmaster can keep a reputation of being very fair, while in fact saving real money for his company. And in 2/3\nicefrac{{2}}{{3}} of the games with guests that play stay, it will appear in the end that switching would have been the better choice. This should motivate enough of the people studying the show to use the switching strategy when they happen to play. As those switchers will never win, they contribute to the benefit of the company.

Is this possibility realistic? Well, a good showmaster worth his money should be able to read his guest’s mind. Before the game starts he interacted enough to get a good guess on the preferred strategy of the guest. The showmaster might even manage to manipulate the guest into adopting a certain strategy. Using this against his guest can make him look more fair than he is in reality. And if the showmaster is unsure, he just plays evil, to be on the safe side.

The acting guest

There is also a counter strategy available to the guest against the mind-reading showmaster. If the guest manages to act in a way that the showmaster believes the guest is playing stay, whereas the guest actually will play switch, the winning chance of the guest becomes 2/3\nicefrac{{2}}{{3}}. This, however, won’t work too often, as the showmaster will certainly recognize whenever he is being fooled. If this happens too often, he might return to the randomized strategy. Or, if he learns to identify the acting guests, he just plays evil on them and they go home with only a goat. Thus the acting guest takes a real risk. If caught, he loses for sure.

The real show

Most people will not perform the above mathematical consideration explicitly when faced with the situation of a real game show. Their intuition, however, might take many factors into account that the classical textbook version of the Monty Hall problem ignores. Intuition knows that the showmaster could have acted differently, might be evil, that he wants to create a situation of emotional tension, and probably wants to avoid giving away too many expensive cars. The showmaster will likely be perceived as an experienced veteran of many such shows, being well capable of guessing or manipulating the strategy of individual guests. Even if told that the showmaster will always open another door, intuition won’t put too much belief in this claim. There is just too much circumstantial evidence coded into our intuition of real showmasters, and other predators, to behave differently, usually.

As the Monty Hall show is an information game, it would be very natural for the showmaster to try to erase any information his guest has on the proper door. The existence of such an information-erasing strategy for the showmaster might determine the ordinary guest’s intuition. If it is possible for the show master to erase any knowledge, he probably does so and therefore a 1/2\nicefrac{{1}}{{2}} chance to the car being behind their door is better assumed.

It should be said that psychology offers also another very plausible explanation for the assignment of a 1/2\nicefrac{{1}}{{2}} chance. It is well known that people use logical shortcuts in their reasoning, which should also lead to the same result. In psychological investigations of the Monty Hall problem the “insensitivity to prior probability outcomes”, considered by [5], is often mentioned as an explanation of the guest’s behavior. It states that prior probabilities are ignored by humans when given new evidence, even if the new information is worthless. Transferring this to the Monty Hall problem, one can explain why people assign a 1/2\nicefrac{{1}}{{2}} probability of winning for both strategies, switching or staying. Thus, it could well be a mere coincidence that people’s intuition provides here a good guess for the chances.

Still, the preference that people show for staying is not explained by this. Staying is better in case the show master is more evil, mind-reading, or manipulative, as it guarantees a winning chance of 1/3\nicefrac{{1}}{{3}}, whereas switching risks to always loose. The only safe strategy against the maneuvers of the showmaster is to stay with the initial choice, thereby ensuring that one will bring home a car in 1/3\nicefrac{{1}}{{3}} of the cases.

The preference for staying might therefore be explained by the distrust in the intentions of the show master. The opening of the second door is easily perceived as an intentional reaction of the show master to the guest’s move to pick a door. Psychological studies by [6] showed that people pay special attention to whether effects are caused by humans or non-humans. Humans have intentions and these can be bad. As the human brain is evolutionary trained to watch out for deception, the more risky strategy to switch is disfavored, despite the 1/2\nicefrac{{1}}{{2}} chance assignment people make.

Thus, the apparently irrational strategy of staying turns out to be in fact very rational, as it is immune against all the mean tricks the showmaster might use. And the apparent contradiction between people assigning a 1/2\nicefrac{{1}}{{2}} chance – for whatever psychological reason – to the car being behind the other door, but strongly insisting on staying with their chosen door is also lifted. The former is an estimate with some uncertainty attached to it. The latter is a decision, taking potential losses into account, which tells us that we could be a sure loser if the showmaster is mentally above us.

Did the real Monty Hall knew about all this? We can let him answer this by himself [7]: "But if he [the showmaster] has the choice whether to allow a switch or not, beware. Caveat emptor. It all depends on his mood.”

Acknowledgement.

We thank Karen Redinger Emmendorfer, Tim Sullivan, Martin Reinecke, Ancla Müller, Sebastian Hutschenreuter, Reimar Leike, and two anonymous reviewers for insightful comments and language corrections.

References